# Alpha-element abundance patterns in star-forming regions of the local Universe

### v1

@@ Abstract
Results. After a careful analysis we conclude that the ICF scheme proposed by Izotov et al. (2006) seems to better reproduce the behaviour of Ne/O, S/O and Ar/O. We find that the distribution of Ne/O ratios in H ii regions shows a large dispersion and no clear trend with O/H, indicating that the different ICF(Ne) schemes are not able to provide correct Ne/O ratios for most of these objects. This is not the case for SFGs that show similar linear relations with slightly positive slopes for the distributions of log(Ne/O) with respect to 12+log(O/H) or 12+log(Ne/H). The origin of this abundance pattern may be a combination of a metallicity-dependent dust depletion of O and ICF effects. The log(S/O) versus 12+log(O/H) distribution is consistent with a constant value especially for HII regions and when we consider both types of objects (SFGs + H ii regions). However, the log(S/O) vs. 12+log(S/H) distribution shows a quite tight linear fit with a positive slope. This relation seems to flatten at 12+log(S/H) $\lesssim$ 6.0. We consider that the observed behaviour of S/O with S/H is compatible with some contribution of S produced by SNe Ia. Finally, the behaviour of log(Ar/O) vs. 12+log(O/H) is very similar for H ii regions and SFGs and seems to be independent of the ionisation degree and the type of ICF(Ar) used, whether based on only the ([Ar iii] lines or on the combination of [Ar iii] and [Ar iv] lines. The linear fit to log(Ar/O) vs. 12+log(O/H) indicates a slight decrease of log(Ar/O) as 12+log(O/H) increases. However, the log(Ar/O) vs. 12+log(Ar/H) relation shows an inverse behaviour, with a small positive slope that could indicate a small contribution of Ar produced by SNe Ia.
@@ 1 Introduction
In recent years, and especially with the advent of the James Webb Space Telescope (JWST), metallicity determinations from the analysis of emission line spectra are beginning to be obtained for high-z SFGs (e.g. Arellano-Córdova et al., 2022; Schaerer et al., 2022; Brinchmann, 2023; Curti et al., 2023; Isobe et al., 2023; Nakajima et al., 2023). Even in some objects, the detection of the faint auroral [O iii] $\lambda$4363 line has permitted to determine the electron temperature, $T_{\rm e}$, and to apply the so-called direct method (Dinerstein, 1990; Peimbert et al., 2017) to obtain more precise determinations of the O abundance (e.g. Arellano-Córdova et al., 2022; Schaerer et al., 2022; Curti et al., 2023; Rhoads et al., 2023; Trump et al., 2023; Laseter et al., 2024). The total abundances of Ne, S and Ar have been determined for several SFGs at $z$ = 4 -- 10, offering the opportunity to study the evolution of the Ne/O, S/O and Ar/O abundance ratios over cosmic time (e.g. Arellano-Córdova et al., 2022; Isobe et al., 2023; Marques-Chaves et al., 2024; Bhattacharya et al., 2024; Stanton et al., 2024). In the case of Ne, Arellano-Córdova et al. (2022, 2024), Isobe et al. (2023) and Stanton et al. (2024) find that the Ne/O ratio –comparing its behaviour with that of the local Universe– does not appear to evolve with redshift. Although, on the other hand, Stanton et al. (2024) find evidence of evolution of the Ar/O ratio with redshift.
@@ 2 Description of the sample of H ii regions and star-forming galaxies
In Table 4 we list the 1386 objects included in the present study, indicating their name, whether they correspond to H ii regions or SFGs –both types correspond exactly to 50% of the objects by a happy coincidence– and the reference of their published spectra.
@@ 3 Physical conditions and ionic abundances
We make use of PyNeb 1.1.18 (Luridiana et al., 2015; Morisset et al., 2020) with the atomic dataset presented in Table 3 and the H i  effective recombination coefficients from Storey & Hummer (1995) to determine the physical conditions and the ionic abundances of the ionised nebulae.
@@ 4 Total O, Ne, S and Ar abundances
As it is said in Sect. 1, we have to consider ICFs to determine the total abundances of Ne, S, and Ar. Following a philosophy similar to that of previous works in H ii regions and SFGs (e.g. Arellano-Córdova et al., 2020, 2024), we have compared the elemental abundances and their ratios obtained using different correction schemes to achieve the most model-independent and robust total abundance determinations possible. For each element, we compare and analyse the results using three ICF schemes. Two of them: Amayo et al. (2021) and Izotov et al. (2006) are common to Ne, S and Ar, and the third scheme is different for each element, Dors et al. (2013) for Ne, Dors et al. (2016) for S and Pérez-Montero et al. (2007) for Ar. In Fig. 14, in the Appendix, we include the differences between the Ne, S and Ar abundances (in logarithmic form) obtained for the spectra of the DESIRED-E sample using the different ICF schemes. The shape of the curves helps to interpret the distribution of abundance ratios that will be discussed below.
@@ 4.1 Neon
Figure 2: Log(Ne2+/O2+) (top) and log((S++S2+)/O) (bottom) as a function of the ionisation degree, O2+/O, for DESIRED-E sample of H ii regions (blue circles) and SFGs (black squares). The black dashed lines and the grey bands show the solar log(Ne/O) and log(S/O) values and their associated uncertainty, respectively, from Asplund et al. (2021). The different curves represent the ICF schemes used in this study: Amayo et al. (2021, red dashed-dotted lines), Izotov et al. (2006, dotted lines with colors corresponding to three metallicity ranges), Dors et al. (2013, green dashed line in top panel) and Dors et al. (2016, green dashed line in bottom panel).
@@
As indicated previously, we use three different ICF(Ne) [^1] Rigorously, in the form we define our ICFs, they should be expressed as ICF(Ne2+/O2+) for Ne, ICF($\frac{{\rm S}^{+}+{\rm S}^{2+}}{\rm O}$) for S, and ICF(Ar2+/O2+) or ICF($\frac{{\rm Ar}^{2+}+{\rm Ar}^{3+}}{{\rm O}^{2+}}$) for the two different cases of Ar. To make the text easier to read, we decided to abbreviate them simply as ICF(Ne), ICF(S) and ICF(Ar). schemes to derive the total Ne abundance. The set by Amayo et al. (2021) is based on a grid of photoionisation models from the Mexican Million Models data base (3MdB Morisset, 2009) under the ‘BOND’ reference (Vale Asari et al., 2016). This grid was selected by applying several filters to resemble the properties of a large observational sample of extragalactic H II regions and the integrated spectra of SFGs. For their ionisation models, Amayo et al. (2021) adopt stellar spectral energy distributions (SEDs) obtained with the PopStar code (Mollá et al., 2009). All the ICFs of Amayo et al. (2021) used in this study depend on the mean ionisation degree of the nebulae –parameterized by the O2+/O ratio– and are, in principle, valid for the whole range of ionisation degrees. The second ICF(Ne) scheme we have used is that by Izotov et al. (2006), who use the grid of photoionisation models of SFGs presented by Stasińska & Izotov (2003) and input stellar SED models from Starburst99 (Leitherer et al., 1999). The ICFs by Izotov et al. (2006) depend on the ionisation degree and the metallicity. They propose different equations for ’low’ (12+log(O/H) $\leq$ 7.2), ’intermediate’ (7.2 $<$ 12+log(O/H) $<$ 8.2) and ’high’ (12+log(O/H) $\geq$ 8.2) metallicity that are used to interpolate the appropriate value of the ICF using the O/H ratio of the object. In the case of the ICF(Ne), the relations given by Izotov et al. (2006) are valid for any value of the O2+/O ratio. The last ICF(Ne) scheme considered in this study is that by Dors et al. (2013), which is based on photoionisation models of extragalactic H ii regions and SFGs that use ionising SEDs from Starburst99 (Leitherer et al., 1999). They provide an analytical expression (their equation 14, applicable for data obtained from optical spectra) parameterized by the O2+/O ratio and valid for all the range of ionisation degrees.
@@
In the top panel of Fig. 2, we plot log(Ne2+/O2+) as a function of O2+/O for 557 spectra of H ii regions and 671 of SFGs of the DESIRED-E sample (1228 in total). The coloured curves represent the different ICF(Ne) values defined by Eq. 1 provided by each of the ICF schemes considered: Amayo et al. (2021, red dashed-dotted lines), Izotov et al. (2006, dotted lines with colors corresponding to the three metallicity ranges) and Dors et al. (2013, green dashed line). The horizontal black dashed line indicates the solar log(Ne/O) of $-0.63\pm 0.06$ –along as its uncertainty represented by the grey band– as reference Asplund et al. (2021), this line also represents the classical ICF of Peimbert & Costero (1969), that assumes Ne2+/O2+ = Ne/O. Fig. 2 clearly indicates that the dispersion of the log(Ne2+/O2+) values of objects classified as H ii regions is much larger than of those classified as SFGs. This observational fact was reported by Kennicutt et al. (2003) and Croxall et al. (2016), among others. Fig. 2 also shows that the ICF(Ne) by Amayo et al. (2021) seems to better reproduce the behaviour of log(Ne2+/O2+) for objects with O2+/O $<$ 0.5 (mostly H ii regions), but giving fairly higher values of log(Ne/O) –between 0.1 and 0.3 dex higher– than the ICF(Ne) schemes by Izotov et al. (2006) and Dors et al. (2013) in the whole ionisation range. This can be noted very clearly in the different distribution of the blue and magenta points in the top left panel of Fig. 14, where we compare the abundance values obtained from the different ICF schemes as a function of the O2+/O ratio.
@@
Fig. 3 illustrates that all the ICF(Ne) schemes considered fail to give correct Ne/O ratios for nebulae having O2+/O $\leq$ 0.5, that would correspond mainly to intermediate and high-metallicity objects. The vast majority of the objects of the DESIRED-E sample in that part of the log(Ne/O) vs. O2+/O diagram correspond to H ii regions. In fact, this is an expected result according to the behaviour of the ICF(Ne) curves against the observed distribution of the Ne2+/O2+ ratios of the DESIRED-E objects shown in the upper panel of Fig. 2.
@@ 4.2 Sulfur
In the bottom panel of Fig. 2, we plot ${}^{+}+$S2+)/O) as a function of the ionisation degree, O2+/O, for 509 H ii regions and 493 SFGs of the DESIRED-E sample with measurements of both, [S ii] and [S iii] lines. The coloured curves represent the behaviour of the different ICF(S) schemes considered: Amayo et al. (2021, red dashed-dotted lines), Izotov et al. (2006, dotted lines with colors corresponding to the three metallicity ranges) and Dors et al. (2016, green dashed line). The horizontal black dashed line shows the value of the solar log(S/O) ratio –along as its uncertainty represented by the grey band– recommended by Asplund et al. (2021). Fig. 2 shows that from values of O2+/O larger than about 0.5, the contribution of S3+ in the nebulae (with an IP of 34.8 eV) becomes important. In the figure, we can see that only the ICF(S) scheme by Izotov et al. (2006) reproduces the observed ${}^{+}+$S2+)/O) values for very high ionisation degree objects (with O2+/O $\geq$ 0.9), mostly corresponding to SFGs.
@@
The two panels of the middle row of Fig. 4 show the results using the ICF(S) by Izotov et al. (2006), which clearly provides flatter trends than the other schemes. In fact, the maximum variation between the extreme values of the log(S/O) of the bins represented in both panels is only of the order of 0.1 dex. Another argument to claim that the ICF(S) by Izotov et al. (2006) is more appropriate is the shortage of objects with very low log(S/O) at O2+/O $\geq$ 0.95 –explained by the curves represented in Figs. 2 and 14– which appear when the other two IFC(S) are applied. The mean log(S/O) obtained with the different ICF(S) schemes are fairly similar: --1.62 $\pm$ 0.19 for Amayo et al. (2021), --1.59 $\pm$ 0.17 for Izotov et al. (2006) and --1.60 $\pm$ 0.19 for Dors et al. (2016). All of these values are consistent with the solar value of --1.57 $\pm$ 0.05 (Asplund et al., 2021). In Fig. 14 we can see that the three ICF(S) used give very similar log(S/H) values when the objects are in the interval 0.1$<$O2+/O$<$0.5, in that case the differences are of the order or even smaller than 0.05 dex.
Figure 5: Log(Ar2+/O2+) (top) and log((Ar2++Ar3+)/O2+) (bottom) as a function of the ionisation degree, O2+/O, for DESIRED-E sample of H ii regions (blue circles) and SFGs (black squares). The black dashed lines and the grey bands show the solar log(Ar/O) value and their associated uncertainty, respectively, from Asplund et al. (2021). The different curves represent the ICF(Ar) schemes used in this study: Amayo et al. (2021, red dashed-dotted lines), Izotov et al. (2006, dotted lines with colors corresponding to three metallicity ranges) and Pérez-Montero et al. (2007, green dashed lines).

### v2

@@ Abstract
Results. After a careful analysis we conclude that one of the tested ICF schemes provides a better match to the observed behaviour of Ne/O, S/O and Ar/O ratios. We find that the distribution of Ne/O ratios in H ii regions shows a large dispersion and no clear trend with O/H, indicating that the different ICF(Ne) schemes are not able to provide correct Ne/O ratios for most of these objects. This is not the case for SFGs that show similar linear relations with slightly positive slopes for the distributions of log(Ne/O) with respect to 12+log(O/H) or 12+log(Ne/H). The origin of this abundance pattern may be a combination of a metallicity-dependent dust depletion of O and ICF effects. The log(S/O) versus 12+log(O/H) distribution is consistent with a constant value especially for HII regions and when we consider both types of objects (SFGs + H ii regions). However, the log(S/O) vs. 12+log(S/H) distribution shows a quite tight linear fit with a positive slope. This relation seems to flatten at 12+log(S/H) $\lesssim$ 6.0. We consider that the observed behaviour of S/O with S/H is compatible with some contribution of S produced by SNe Ia. Finally, the behaviour of log(Ar/O) vs. 12+log(O/H) is very similar for H ii regions and SFGs and seems to be independent of the ionisation degree and the type of ICF(Ar) used, whether based on only the ([Ar iii] lines or on the combination of [Ar iii] and [Ar iv] lines. The linear fit to log(Ar/O) vs. 12+log(O/H) indicates a slight decrease of log(Ar/O) as 12+log(O/H) increases. However, the log(Ar/O) vs. 12+log(Ar/H) relation shows an inverse behaviour, with a small positive slope that could indicate a small contribution of Ar produced by SNe Ia.
@@ 1 Introduction
In recent years, and especially with the advent of the James Webb Space Telescope (JWST), metallicity determinations from the analysis of emission line spectra are beginning to be obtained for high-z SFGs (e.g. Arellano-Córdova et al., 2022; Schaerer et al., 2022; Brinchmann, 2023; Curti et al., 2023; Isobe et al., 2023; Nakajima et al., 2023). Even in some objects, the detection of the faint auroral [O iii] $\lambda$4363 line has permitted to determine the electron temperature, $T_{\rm e}$, and to apply the so-called direct method (Dinerstein, 1990; Peimbert et al., 2017) to obtain more precise determinations of the O abundance (e.g. Arellano-Córdova et al., 2022; Schaerer et al., 2022; Curti et al., 2023; Rhoads et al., 2023; Trump et al., 2023; Laseter et al., 2024). The total abundances of Ne, S and Ar have been determined for several SFGs at $z$ = 4 -- 10, offering the opportunity to study the evolution of the Ne/O, S/O and Ar/O abundance ratios over cosmic time (e.g. Arellano-Córdova et al., 2022; Isobe et al., 2023; Marques-Chaves et al., 2024; Stanton et al., 2025). In the case of Ne, Arellano-Córdova et al. (2022, 2024), Isobe et al. (2023) and Stanton et al. (2025) find that the Ne/O ratio –comparing its behaviour with that of the local Universe– does not appear to evolve with redshift. Although, on the other hand, Stanton et al. (2025) find evidence of evolution of the Ar/O ratio with redshift.
@@ 2 Description of the sample of H ii regions and star-forming galaxies
In Table D.2 we list the 1386 objects included in the present study, indicating their name, whether they correspond to H ii regions or SFGs –both types correspond exactly to 50% of the objects by a happy coincidence– and the reference of their published spectra.
@@ 3 Physical conditions and ionic abundances
We make use of PyNeb 1.1.18 (Luridiana et al., 2015; Morisset et al., 2020) [^1] PyNeb code is publicly available on GitHub: https://github.com/Morisset/PyNeb_devel. with the atomic dataset presented in Table D.1 and the H i  effective recombination coefficients from Storey & Hummer (1995) to determine the physical conditions and the ionic abundances of the ionised nebulae. PyNeb is written in Python programming language and is fully vectorized. To calculate the emissivity of collisional excitation lines, PyNeb calculates the relative population of the atomic levels of the ions of interest by solving the statistical equilibrium equations. In the case of recombination lines, it interpolates the available emissivity tables.
@@ 4 Total O, Ne, S and Ar abundances
As it is said in Sect. 1, we have to consider ICFs to determine the total abundances of Ne, S, and Ar. Following a philosophy similar to that of previous works in H ii regions and SFGs (e.g. Arellano-Córdova et al., 2020, 2024), we have compared the elemental abundances and their ratios obtained using different correction schemes to achieve the most model-independent and robust total abundance determinations possible. For each element, we compare and analyse the results using three ICF schemes. Two of them: Amayo et al. (2021) and Izotov et al. (2006) are common to Ne, S and Ar, and the third scheme is different for each element, Dors et al. (2013) for Ne, Dors et al. (2016) for S and Pérez-Montero et al. (2007) for Ar. In Fig. 14, in the Appendix, we include the differences between the Ne, S and Ar abundances (in logarithmic form) obtained for the spectra of the DESIRED-E sample using the different ICF schemes. The shape of the curves helps to interpret the distribution of abundance ratios that will be discussed below. Moreover, The range of values covered by the observational points in these figures –which depends on the element, the emission lines used and the degree of ionisation– gives us an idea of the uncertainty introduced by the use of an ICF on the value of the total abundance of a given element and object.
@@ 4.1 Neon
Figure 2: Log(Ne2+/O2+) (top) and log((S++S2+)/O) (bottom) as a function of the ionisation degree, O2+/O, for DESIRED-E sample of H ii regions (blue circles) and SFGs (black squares). The black dashed lines and the grey bands show the solar log(Ne/O) and log(S/O) values and their associated uncertainty, respectively, from Asplund et al. (2021). The different curves represent the ICF schemes used in this study: Amayo et al. (2021, red dashed-dotted lines), Izotov et al. (2006, dotted lines with colours corresponding to three metallicity ranges), Dors et al. (2013, green dashed line in top panel) and Dors et al. (2016, green dashed line in bottom panel).
@@
As indicated previously, we use three different ICF(Ne) [^2] Rigorously, in the form we define our ICFs, they should be expressed as ICF(Ne2+/O2+) for Ne, ICF($\frac{{\rm S}^{+}+{\rm S}^{2+}}{\rm O}$) for S, and ICF(Ar2+/O2+) or ICF($\frac{{\rm Ar}^{2+}+{\rm Ar}^{3+}}{{\rm O}^{2+}}$) for the two different cases used for Ar. To make the text easier to read, we decided to abbreviate them simply as ICF(Ne), ICF(S) and ICF(Ar). schemes to derive the total Ne abundance. The set by Amayo et al. (2021) is based on a grid of photoionisation models from the Mexican Million Models data base (3MdB Morisset, 2009) under the ‘BOND’ reference (Vale Asari et al., 2016). This grid was selected by applying several filters to resemble the properties of a large observational sample of extragalactic H II regions and the integrated spectra of SFGs. For their ionisation models, Amayo et al. (2021) adopt stellar spectral energy distributions (SEDs) obtained with the PopStar code (Mollá et al., 2009). All the ICFs of Amayo et al. (2021) used in this study depend on the mean ionisation degree of the nebulae –parameterized by the O2+/O ratio– and are, in principle, valid for the whole range of ionisation degrees. The second ICF(Ne) scheme we have used is that by Izotov et al. (2006), who use the grid of photoionisation models of SFGs presented by Stasińska & Izotov (2003) and input stellar SED models from Starburst99 (Leitherer et al., 1999). The ICFs by Izotov et al. (2006) depend on the ionisation degree and the metallicity. They propose different equations for ’low’ (12+log(O/H) $\leq$ 7.2), ’intermediate’ (7.2 $<$ 12+log(O/H) $<$ 8.2) and ’high’ (12+log(O/H) $\geq$ 8.2) metallicity that are used to interpolate the appropriate value of the ICF using the O/H ratio of the object. In the case of the ICF(Ne), the relations given by Izotov et al. (2006) are valid for any value of the O2+/O ratio. The last ICF(Ne) scheme considered in this study is that by Dors et al. (2013), which is based on photoionisation models of extragalactic H ii regions and SFGs that use ionising SEDs from Starburst99 (Leitherer et al., 1999). They provide an analytical expression (their equation 14, applicable for data obtained from optical spectra) parameterized by the O2+/O ratio and valid for all the range of ionisation degrees.
@@
In the top panel of Fig. 2, we plot log(Ne2+/O2+) as a function of O2+/O for 557 spectra of H ii regions and 671 of SFGs of the DESIRED-E sample (1228 in total). The coloured curves represent the different ICF(Ne) values defined by Eq. 1 provided by each of the ICF schemes considered: Amayo et al. (2021, red dashed-dotted lines), Izotov et al. (2006, dotted lines with colours corresponding to the three metallicity ranges) and Dors et al. (2013, green dashed line). The horizontal black dashed line indicates the solar log(Ne/O) of $-0.63\pm 0.06$ –along as its uncertainty represented by the grey band– as reference Asplund et al. (2021), this line also represents the classical ICF of Peimbert & Costero (1969), that assumes Ne2+/O2+ = Ne/O. Fig. 2 clearly indicates that the dispersion of the log(Ne2+/O2+) values of objects classified as H ii regions is much larger than of those classified as SFGs. This observational fact was reported by Kennicutt et al. (2003) and Croxall et al. (2016), among others. Fig. 2 also shows that the ICF(Ne) by Amayo et al. (2021) seems to better reproduce the behaviour of log(Ne2+/O2+) for objects with O2+/O $<$ 0.5 (mostly H ii regions), but giving fairly higher values of log(Ne/O) –between 0.1 and 0.3 dex higher– than the ICF(Ne) schemes by Izotov et al. (2006) and Dors et al. (2013) in the whole ionisation range. This can be noted very clearly in the different distribution of the blue and magenta points in the top left panel of Fig. 14. The Ne abundances obtained using the different ICFs can differ by up to 0.3 dex when O2+/O $<$ 0.5, although this difference decreases at higher ionisation degrees, reaching values below 0.1 dex when O2+/O $>$ 0.8.
@@
Fig. 3 illustrates that all the ICF(Ne) schemes considered fail to give correct Ne/O ratios for nebulae having O2+/O $\leq$ 0.5, that would correspond mainly to intermediate and high-metallicity objects. The vast majority of the objects of the DESIRED-E sample in that part of the log(Ne/O) vs. O2+/O diagram correspond to H ii regions. In fact, as this is an expected result according to the behaviour of the ICF(Ne) curves against the observed distribution of the Ne2+/O2+ ratios of the DESIRED-E objects shown in the upper panel of Fig. 2. The total O and Ne abundances determined using the different ICF schemes considered –as well as the total abundances of S and Ar– are collected in Table D5.
@@ 4.2 Sulfur
In the bottom panel of Fig. 2, we plot ${}^{+}+$S2+)/O) as a function of the ionisation degree, O2+/O, for 509 H ii regions and 493 SFGs of the DESIRED-E sample with measurements of both, [S ii] and [S iii] lines. The coloured curves represent the behaviour of the different ICF(S) schemes considered: Amayo et al. (2021, red dashed-dotted lines), Izotov et al. (2006, dotted lines with colours corresponding to the three metallicity ranges) and Dors et al. (2016, green dashed line). The horizontal black dashed line shows the value of the solar log(S/O) ratio –along as its uncertainty represented by the grey band– recommended by Asplund et al. (2021). Fig. 2 shows that from values of O2+/O larger than about 0.5, the contribution of S3+ in the nebulae (with an IP of 34.8 eV) becomes important. In the figure, we can see that only the ICF(S) scheme by Izotov et al. (2006) reproduces the observed ${}^{+}+$S2+)/O) values for very high ionisation degree objects (with O2+/O $\geq$ 0.9), mostly corresponding to SFGs.
@@
The two panels of the middle row of Fig. 4 show the results using the ICF(S) by Izotov et al. (2006), which clearly provides flatter trends than the other schemes. In fact, the maximum variation between the extreme values of the log(S/O) of the bins represented in both panels is only of the order of 0.1 dex. Another argument to claim that the ICF(S) by Izotov et al. (2006) is more appropriate is the shortage of objects with very low log(S/O) at O2+/O $\geq$ 0.95, which appears when the other two IFC(S) are applied and which is explained by the sharp drop in the curves at those high ionisation degrees that can be seen in Figs. 2 and 14. The mean log(S/O) obtained with the different ICF(S) schemes are fairly similar: --1.62 $\pm$ 0.19 for Amayo et al. (2021), --1.59 $\pm$ 0.17 for Izotov et al. (2006) and --1.60 $\pm$ 0.19 for Dors et al. (2016). All of these values are consistent with the solar value of --1.57 $\pm$ 0.05 (Asplund et al., 2021). In Fig. 14 we can see that the three ICF(S) used give very similar log(S/H) values when the objects are in the interval 0.1$<$O2+/O$<$0.5, in that case the differences are of the order or even smaller than 0.05 dex.
Figure 5: Log(Ar2+/O2+) (top) and log((Ar2++Ar3+)/O2+) (bottom) as a function of the ionisation degree, O2+/O, for DESIRED-E sample of H ii regions (blue circles) and SFGs (black squares). The black dashed lines and the grey bands show the solar log(Ar/O) value and their associated uncertainty, respectively, from Asplund et al. (2021). The different curves represent the ICF(Ar) schemes used in this study: Amayo et al. (2021, red dashed-dotted lines), Izotov et al. (2006, dotted lines with colours corresponding to three metallicity ranges) and Pérez-Montero et al. (2007, green dashed lines).